Grain worth 65 cents per pound was mixed with grain worth 90 cents per pound to produce 20 pounds of a mixture worth 80 cents per pound. Find how many pounds of each was used
Answer by checkley77(12844)

65x+90(20-x)=80*20
65x+1800-90x=1600
-25x=1600-1800
-25x=-200
x=-200/-25
x=8 pounds of 65 cent grain.
20-8=12 pounds of 90 cent grain.
Proof:
65*8+90*12=1600
520+1080=1600
1600=1600